OpenStudy (isharp):

What are the zeroes of f(x) = x2 − 6x + 8? A.) x = −4, 2 B.) x = −4, −2 C.) x = 4, 2 D.)x = 4, −2 whats the answer

OpenStudy (sheriph05):

Find two numbers when you add/subtract gives -6 and when you multiply gives positive 8. Replace -6 with the two numbers found, and factor out to obtain two values of x. The value you obtain is the answer...You can try and let me know if you encounter any problem.

OpenStudy (mrs.ambrose614):

The awnser is x=4,2 I hope I helped you set the function equal to 0 and solve for x @isharp

jhonyy9 (jhonyy9):

f(x) = x^2 -6x +8 = (x-4)(x-2) factorized completely x-4=0 x_1=4 and x-2=0 x_2=2 so the roots are 4 and 2 so choice c. hope helped